Find the slope of the line passing through the points (4,8) and (8,7).
Answer by rfer(16322)

slope m
m=(7-8)/(8-4)
m=-1/4
----------------
y=mx+b
7=-1/4*8+b
7=-2+b
7+2=b
b=9
------------
y=-1/4x+9